Lorge Corporation has collected the following information after its first year of sales. Sales were $1,600,000 on 80,000 units; selling expenses $250,000 (40% variable and 60% fixed); direct materials $620,800; direct labor $270,000; administrative expenses $270,000 (20% variable and 80% fixed); and manufacturing overhead $336,000 (70% variable and 30% fixed). Top management has asked you to do a CVP analysis so that it can make plans for the coming year. It has projected that unit sales will increase by 10% next year.

Compute (1) the contribution margin for the current year and the projected year, and (2) the fixed costs for the current year. (Assume that fixed costs will remain the same in the projected year.)

(1) Contribution margin for current year

$

Contribution margin for projected year

$

(2) Fixed costs for current year

$

eTextbook and Media

Compute the break-even point in units and sales dollars for the first year. (Round contribution margin ratio to 2 decimal places e.g. 0.15 and final answers to 0 decimal places, e.g. 2,510.)

Break-even point units
Break-even point

$

eTextbook and Media

The company has a target net income of $190,000. What is the required sales in dollars for the company to meet its target?

Sales dollars required for target net income

$

eTextbook and Media

If the company meets its target net income number, by what percentage could its sales fall before it is operating at a loss? That is, what is its margin of safety ratio? (Round answer to 1 decimal place, e.g. 10.5.)

Margin of safety ratio %

eTextbook and Media

The company is considering a purchase of equipment that would reduce its direct labor costs by $110,000 and would change its manufacturing overhead costs to 30% variable and 70% fixed (assume total manufacturing overhead cost is $336,000, as above). It is also considering switching to a pure commission basis for its sales staff. This would change selling expenses to 90% variable and 10% fixed (assume total selling expense is $250,000, as above). Compute (1) the contribution margin and (2) the contribution margin ratio, and recompute (3) the break-even point in sales dollars. (Round contribution margin ratio to 2 decimal places, e.g. 25.25 and all other answers to 0 decimal places, e.g. 2,520. Use the current year numbers for calculations.)

1. Contribution margin

$

2. Contribution margin ratio %
3. Break-even point

$
Answer #1
Sales $    16,00,000
Less Variable Expenses
Direct Material $      6,20,800
Direct Labor $      2,70,000
Administrative expenses $          54,000 =270000*20%
Manufacturing Overhead $      2,35,200 =336000*70%
Selling Expenses $      1,00,000 =250000*40%
Total Variable expenses $    12,80,000
Contribution Margin $      3,20,000
Less Fixed Costs
Administrative expenses $      2,16,000 =270000*80%
Manufacturing Overhead $      1,00,800 =336000*30%
Selling Expenses $      1,50,000 =250000*60%
Total Fixed Expenses $      4,66,800
Net Operating Income $     -1,46,800


Contribution Margin for Current Year = $320000
Contribution Margin for Projected Year = $320000 x 1.1 = $352000
Fixed Expenses = $466800

Contribution Margin per unit = $320000 / 80000 = $4 per unit
Contribution Margin ratio = $4/$20 x 100 = 20%

Break Even Point (Units) = Fixed Expenses / Contribution margin per unit
= $466800 / $4 = 116700 units

Break Even Point ($) = Fixed Expenses / Contribution margin Ratio
= $466800 / 20% = $2334000

Sales dollars required for target net income = (Fixed Expenses+Desired Profit) / Contribution margin Ratio
= ($466800+190000) / 20% = $3284000

Margin of Safety = $3284000 - $2334000 = $950000
Margin of Safety (%) = $950000 / $3284000 x 100 = 28.93%

Q2.

Sales $    16,00,000
Less Variable Expenses
Direct Material $      6,20,800
Direct Labor $      1,60,000 =270000-110000
Administrative expenses $          54,000 =270000*20%
Manufacturing Overhead $      1,00,800 =336000*30%
Selling Expenses $      2,25,000 =250000*90%
Total Variable expenses $    11,60,600
Contribution Margin $      4,39,400
Less Fixed Costs
Administrative expenses $      2,16,000 =270000*80%
Manufacturing Overhead $      2,35,200 =336000*70%
Selling Expenses $          25,000 =250000*10%
Total Fixed Expenses $      4,76,200
Net Operating Income $        -36,800


Contribution Margin = $439400
Contribution Margin ratio = $439400 / $1600000 x 100 = 27.4625%
Break Even Point = $476200 / 27.4625% = $1734001